# Deep learning for classifying dynamical states from time series via recurrence plots

### v1

@@ Abstract
Recurrence Quantification Analysis (RQA) is a widely used method for capturing the dynamical structure embedded in time series data, relying on the analysis of recurrence patterns in the reconstructed phase space via recurrence plots. Although RQA proves effective across a range of applications, it typically requires the computation of multiple quantitative measures, making it both computationally intensive and sensitive to parameter choices. In this study, we adopt an alternative approach that bypasses manual feature selection and extraction by directly using recurrence plot images as input to a deep learning model. We propose a new dual-branch deep learning architecture specifically designed to efficiently capture the complex dynamical features encoded in RPs. We also compare its performance against a baseline ResNet-50 model for classifying the dynamical behavior of time series using recurrence plots. Our dual-branch model, trained exclusively on simulated time series, accurately and efficiently distinguishes among six distinct classes: periodic, quasi-periodic, chaotic, hyperchaotic, white noise, and red noise. To assess its generalizability, we apply the trained model to time series generated from standard Lorenz and Rössler systems, neither of which is included in the training set, as well as to experimental datasets from a Chua circuit and observational light curves of the variable stars AC Her, SX Her, and Chi Cygni. In all cases, the model outperforms the baseline and yields predictions that align with the known dynamics of these systems. These results further demonstrate the robustness and versatility of our deep learning framework, underscoring the potential of RP based models as fast, accurate, and scalable tools for classifying dynamical states in both synthetic and real-world time series data.
I Introduction
@@
Understanding and classifying emergent dynamical states in real-world and engineered systems remains a fundamental challenge due to the intrinsic complexity of their physical mechanisms and the involvement of multiple interacting parameters. However, the increasing availability of large scale time series data and advances in high performance computing open new opportunities to uncover meaningful dynamical patterns by integrating data driven approaches with computational modeling. Many real systems such as biological, climate, and social systems and engineered systems including mechanical, electronics and power grid systems often generate nonlinear time series that encode a variety of rich dynamical behaviors, including periodic, quasiperiodic, chaotic, and hyperchaotic regimes [1, 2, 3, 4, 5]. Accurate identification of these behaviors is essential not only for advancing our understanding of the underlying physics but also for anticipating critical transitions, detecting early warning signals, and designing robust prediction and control strategies [6, 7, 8, 9].
@@
Recurrence quantification analysis (RQA) has emerged as one of the most powerful techniques for uncovering dynamical states from time series data, whether derived from dynamical models or empirical measurements. RQA has been successfully applied across a wide range of disciplines, for example, in biological systems for the analysis of heart rate variability analysis [13] and neural activity characterization [14]; in climate science to identify climate change dynamics [15, 16, 17]; and in engineering systems for the prediction of mechanical failures [18]. This method analyzes recurrence patterns in the reconstructed phase space from data using recurrence plots (RPs). Introduced by Eckmann et al. [19], RPs provide a visual representation of the times when the state of the system returns close to the states visited previously[20]. They reveal complex dynamical features, such as periodicity, laminar phases, drifts, and chaotic bursts, that often escape detection by traditional linear or spectral methods [21, 22]. RQA extends the basic recurrence plot framework by introducing a set of quantitative measures, such as recurrence rate, determinism, laminarity, and entropy, that enable systematic analysis of the underlying dynamics [23]. However, as demonstrated in a recent study [24], these measures often exhibit overlapping value ranges across different dynamical regimes, making it difficult to uniquely identify a specific state based on any single measure. Reliable classification, therefore, typically requires a carefully chosen combination of multiple RQA features, underscoring the limitations of relying on individual measures for accurate identification of dynamical states.
@@
A promising alternative bypasses the need for explicit feature extraction altogether by feeding recurrence plots directly into deep learning models, particularly convolutional neural networks (CNNs). CNNs, widely used in image classification, automatically learn hierarchical spatial features through convolutional layers. This image based strategy enables CNNs to extract discriminative patterns directly from recurrence plots without requiring prior feature engineering. Notable applications include arrhythmia detection from ECG data [25], activity recognition using wearable sensors [28], control system development [29], and parameter estimation in nonlinear systems [30]. These studies show that RP based deep learning models can match or even outperform traditional feature based approaches, offering greater scalability and adaptability. Despite these advances, CNN architectures suffer from vanishing gradients, which hinder the training of deeper models. This challenge is largely resolved with the introduction of ResNet architectures [31], which use residual connections to preserve gradient flow and enable deep network training.
In our study, we propose a deep learning model that incorporates ResNet-50[31] as a backbone for feature extraction from recurrence plots, followed by two parallel branches inspired by the DSRNet architecture [32]. These branches are designed to emphasize features corresponding to various directional edges, which are important for recurrence plots. We then compare the performance of our proposed model against a standard ResNet-50 implementation. In the proposed deep learning model, we provide RPs as input images, and train the model to classify time series into one of the six categories: periodic, quasi-periodic, chaotic, hyperchaotic, white noise, or red noise. We demonstrate that our approach offers higher accuracy and greater computational efficiency compared to traditional RQA-based methods and other CNN-based methods that uses RPs as input images.
@@
Our paper is organized as follows. In section II we discuss the methodology for the classification of nonlinear time series by integrating machine learning models and recurrence plots. Section III presents the results that show the performance of our proposed model for simulation and real-world data sets. In the concluding section  IV, we summarize the results and indicate promising future extensions for our model.
II Methodology
@@
To construct the training dataset, we first simulate a set of standard dynamical systems along with noisy data sets, as summarized in Table 1. From the resulting time series, we reconstruct the phase space using Takens’ embedding theorem. Recurrence plots are then generated from these reconstructed trajectories to capture the intrinsic recurrence structure of the underlying dynamics.
@@
These RPs are used as input features for the proposed deep learning framework, which is designed to classify distinct dynamical regimes and noise types. A schematic representation of the overall methodology that includes a systematic way of classification of time series using deep learning model is shown in Figure 1. Comprehensive details on time series generation, RP construction, and the architecture of the proposed deep learning model are provided in the following subsections.
II.1 Time series datasets
@@
To generate the time series datasets, we simulate standard nonlinear dynamical systems as listed in Table 1 using the RK45 integrator. By systematically varying the model parameters and initial conditions, we ensure a broad coverage of nonlinear behaviors. For instance, Chen 4D and Lorenz 4D systems are used to generate time series exhibiting chaotic, hyperchaotic, and periodic dynamics. We use the third system to generate time series that show quasi-periodic behavior. For each simulated system, we use a single time series corresponding to the $x$ variable (see Fig. 1(a)). In addition, we generate noisy signals, of Gaussian white noise and red noise, using the algorithm descirbed by Timmer and Koenig[33]. We also produce additional noisy datasets by adding Gaussian noise to the state variables of the deterministic systems. The standard deviation of the noise is set as a fixed proportion of the original data, as described below:
@@
$\text{with}\quad\sigma_{n}$ $=\alpha\cdot\sigma_{x},\quad\alpha\sim\mathcal{U}(0,0.05),$
where $\sigma_{x}$ denotes the standard deviation of the data $x(t)$ obtained from the dynamical system, and $\sigma_{n}$ denotes the standard deviation of noise. We normalize each dataset before using it to reconstruct the phase space and analyze recurrence features, as detailed in the following section.
Table 1: Dynamical and stochastic systems used to generate different classes of time series.
@@
Chen 4D[34] $\dfrac{dx}{dt}=a(y-x)+e\,yz$ $\dfrac{dy}{dt}=cx-d\,xz+y+u$ $\dfrac{dz}{dt}=xy-bz$ $\dfrac{dw}{dt}=-k\,y$ $a=35,\;b=4.9,\;c=25,\;d=5,\;e=35$ $k=100$ $k=315$ $k=495$ Hyperchaotic Chaotic Periodic
[…]
Quasiperiodic System[36] $\dfrac{dx}{dt}=y$ $\dfrac{dy}{dt}=(\lambda+z+x^{2}-\beta x^{4})y-\omega_{0}^{2}x$ $\dfrac{dz}{dt}=b(\varepsilon-z)-ky^{2}$ $\lambda=-1,\;c=4,\;\beta=\tfrac{1}{18},$ $\omega_{0}=2\pi,\;b=1,\;k=0.02$ Quasiperiodic
Stochastic Data[33] $N_{1},N_{2}\sim\mathcal{N}(0,1)$ $S(f_{k})\sim f_{k}^{-\beta}$ $F(f_{k})=\sqrt{\dfrac{S(f_{k})}{2}}(N_{1}+iN_{2})$ $x(t)=\mathcal{F}^{-1}\left[F(f_{k})\right]$ $\beta=0$ $\beta=2$ White Noise Red Noise
Figure 1: Classification of time series. Schematic representation of the proposed methodology for classifying time series using RPs as input to a deep learning model. (a) We begin by selecting a normalized time series from some simulated dynamical systems and noise generator as listed in Table 1 (b) The phase space of the selected time series is reconstructed using Takens’ embedding theorem. (c) A recurrence plot is then generated from the embedded trajectory using a suitable recurrence threshold $\epsilon$. (d) The resulting RP is resized and used as the input to the deep learning model. In this model, ResNet-50 helps to extract features from RP. Its output is processed through two parallel branches: one capturing spatial features and the other emphasizing directional sensitivity through fixed orientation filters. The outputs from both branches are concatenated and passed through a dense layer. Finally, a softmax layer generates class wise probabilities, from which the most probable dynamical state is identified. For example, if the input time series is chaotic, the output layer assigns a higher probability to the chaotic class.
II.2 Reconstruction of Phase space
From the normalized univariate time series, we reconstruct the system’s phase space using Takens’ embedding theorem [37]. According to this theorem, the topological structure of the original multi dimensional state space can be recovered by constructing delay vectors with appropriate delay and embedding dimension.
@@ II.2 Reconstruction of Phase space
Various techniques are available to estimate appropriate values for the time delay $\tau$ and the embedding dimension $d$, which are essential to reconstruct the trajectory of the system from a single scalar time series. In our study, we employ the mutual information method [38] to estimate $\tau$. The first local minimum of the mutual information curve is typically selected as the optimal delay, as it marks the point at which successive values of the time series become sufficiently independent for effective phase space reconstruction. Although the precise value of the delay may depend on the resolution of the time series, the method remains robust provided that the search is performed over a sufficiently large range to capture the local minimum.
To determine the embedding dimension $d$, we use the method of False Nearest Neighbors (FNN), following the approach proposed by Kennel et al. [39]. The underlying idea is that if the reconstructed trajectory is folded onto itself in a lower dimensional space, it results in a higher number of false neighbors. The knee point at which this number levels off to zero indicates a suitable embedding dimension that preserves the system’s topology.
With the appropriate choice of $\tau$ and $d$, the original univariate time series is transformed into a sequence of embedded vectors. These vectors are organized into a matrix $\mathbf{X}$ as follows:
@@
II.3 Recurrence Plots
Recurrence plots provide a two dimensional representation of the times at which a dynamical system revisits states that are close to previously encountered ones, within a predefined distance threshold $\epsilon$. Given a delay embedded matrix $\mathbf{X}$ in Eq. 3, the recurrence matrix $\mathbf{R}$ is defined as:

### v2

@@ Abstract
Recurrence Quantification Analysis (RQA) is a widely used method for capturing the dynamical structure embedded in time series data, relying on the analysis of recurrence patterns in the reconstructed phase space via recurrence plots (RPs). Although RQA proves effective across a range of applications, it typically requires the computation of multiple quantitative measures, making it both computationally intensive and sensitive to parameter choices. In this study, we adopt an alternative approach that bypasses computation of recurrence measures, by directly using images of RP as input to a deep learning model. We propose a new dual-branch deep learning model named DBResNet-50 built on ResNet-50 architecture and specifically designed to efficiently capture the complex dynamical features encoded in RPs. We compare its performance with standard ResNet-50 and MobileNetV2 and find that DBResNet-50 consistently outperforms both benchmark models. Our DBResNet-50 model, trained exclusively on simulated time series, accurately classifies seven dynamical regimes—periodic, quasi-periodic, chaotic, hyperchaotic, white noise, pink noise, and red noise. Further, to assess its generalizability, we test the trained model on RP images generated from standard dynamical systems not included in the training set, as well as experimental datasets from a Chua circuit, X-ray light curves from the black-hole system GRS 1915+105, and observational light curves of the variable stars AC Her, SX Her, and Chi Cygni. In all cases, DBResNet-50 outperforms the baselines and correctly predicts the known dynamics of these systems. The model further infers the relative contributions of deterministic and stochastic components within a signal, as observed in temperature data from Ladakh and Ranchi. These results demonstrate the robustness and versatility of our deep learning framework and underscore the potential of RP image-based models as fast, accurate, and scalable tools for classifying dynamical states in both synthetic and real-world time series data.
1 Introduction
@@
Understanding and classifying emergent dynamical states in natural and engineered systems remains a fundamental challenge due to the intrinsic complexity of their physical mechanisms and the involvement of multiple interacting parameters. However, the increasing availability of large scale time series data and advances in high-performance computing open new opportunities to uncover meaningful dynamical patterns by integrating data driven approaches with computational modeling. Many real-world systems such as biological, climate, and social systems, and engineered systems, such as mechanical, electronics and power grid systems often generate nonlinear time series that encode a variety of rich dynamical behaviors, including periodic, quasiperiodic, chaotic, and hyperchaotic regimes [19, 35, 10, 37, 26]. Accurate identification of these behaviors is essential not only for advancing our understanding of the underlying physical processes but also for anticipating critical transitions, deriving early warning signals, and designing robust prediction and control strategies [45, 27, 36, 46].
@@
Recently, Recurrence Quantification Analysis (RQA) has emerged as one of the most powerful techniques for uncovering dynamical states from time series data, whether derived from dynamical models or empirical measurements. RQA has been successfully applied across a wide range of disciplines, for example, in biological systems for the analysis of heart rate variability analysis [32] and neural activity characterization [41]; in climate science to identify climate change dynamics [57, 53, 18]; and in engineering systems for the prediction of mechanical failures [39]. This method analyzes recurrence patterns in the reconstructed phase space from data using recurrence plots (RPs). Introduced by Eckmann et al. [11], RPs provide a visual representation of the times when the state of the system returns close to the states visited previously[1]. They reveal complex structure related to dynamical features, such as periodicity, laminar phases, drifts, and chaotic bursts, that often escape detection by traditional linear or spectral methods [31, 23]. The RQA extends the basic recurrence plot framework by introducing a set of quantitative measures, such as recurrence rate, determinism, laminarity, and entropy, that enable systematic analysis of the underlying dynamics [54]. However, as demonstrated in a recent study [50], these measures often exhibit overlapping value ranges across different dynamical regimes, making it difficult to uniquely identify a specific state based on any single measure. Thus, reliable classification, typically requires a carefully chosen combination of multiple RQA features, underscoring the limitations of relying on individual measures for accurate identification of dynamical states.
@@
An alternative approach would be feeding the time series directly into a machine learning model, but they primarily encode statistical correlations and do not reveal the underlying state-space geometry, such as phase space structure or evolution of trajectory explicitly [47]. Moreover, although deep learning models, in principle, can learn internal state representations, they do so in a data-intensive and opaque manner and typically without physical interpretability. In addition, direct use of time series data leads to lower classification accuracy and substantially higher computational cost compared to recurrence based features, as demonstrated in [47]. In contrast, the RPs yield meaningful, and interpretable measures that directly reflect the dynamical properties of the systems. Therefore, we rely on RP images directly to classify dynamical states more accurately and with clearer physical insights.
@@
A promising alternate approach that bypasses the need for explicit computation of RQA mneasures for feature extraction altogether would be feeding recurrence plots directly into deep learning models, particularly convolutional neural networks (CNNs). The CNNs, widely used in image classification, automatically learn hierarchical spatial features through convolutional layers. This image based strategy enables CNNs to extract discriminative patterns directly from recurrence plots without requiring prior feature engineering. A few notable applications in this context include arrhythmia detection from ECG data [34], activity recognition using wearable sensors [13], control system development [12], and parameter estimation in nonlinear systems [30]. These studies show that RP based deep learning models can match or even outperform traditional recurrence measure-based approaches, offering greater scalability and adaptability. Despite these advances, CNN architectures suffer from vanishing gradients, which hinder the training of deeper models. This challenge is largely resolved with the introduction of ResNet architectures [16], which use residual connections to preserve gradient flow and enable deep network training.
In our study, we introduce a dual-branch deep learning model named DBResNet-50, which incorporates ResNet-50[16] as the backbone for feature extraction from images of RPs, followed by two parallel branches inspired by the DSRNet architecture [56]. These branches are designed to emphasize features corresponding to various directional edges, which are highly relevant for RPs. We then evaluate the performance of our proposed model along with the standard ResNet-50. We also consider the MobileNetV2 model [44], which is a lighter architecture with fewer parameters than ResNet-50.
All three models are trained on recurrence plots to classify time series into seven categories—periodic, quasi-periodic, chaotic, hyperchaotic, white noise, pink noise, and red noise, with the noise types chosen due to their prevalence in natural systems. Then, we compare their performances on a variety of data sets, derived from synthetic and real world systems. The results show that our dual-branch architecture achieves the highest accuracy while maintaining superior computational efficiency.
@@
Our paper is organized as follows. In section 2, we discuss the methodology for the classification of nonlinear time series by integrating machine learning models and recurrence plots. Section 3 presents the results that show the performance of the deep learning models for simulation and real-world data sets. In the concluding section  4, we summarize the results and indicate promising future extensions for our model.
2 Methodology
@@
To generate the training and testing datasets, we first simulate a set of standard dynamical systems along with 3 types of noise, as summarized in Table 1. From the resulting time series, we reconstruct the trajectories in phase space using Takens’ embedding theorem and derive the corresponding RPs to capture the intrinsic recurrence structure of the underlying dynamics.
@@
These RPs are then used as input features for the proposed deep learning framework, which is designed to classify distinct dynamical regimes and noise types. A schematic representation of the basic methodology that indicates the systematic way of classification of time series using the dual branch deep learning model is shown in Fig. 1. Comprehensive details on time series generation, RP construction, and the architecture of the proposed deep learning model are provided in the following subsections.
2.1 Time series datasets
@@
To generate the time series datasets, we simulate standard nonlinear dynamical systems as listed in Table 1 using the RK45 integrator. By systematically varying the model parameters and initial conditions, we ensure a broad coverage of dynamical behaviors. For instance, Chen 4D and Lorenz 4D systems are used to generate time series exhibiting chaotic, hyperchaotic, quasi-periodic, and periodic dynamics for choosing appropriate parameters. The Lorenz 3D system is used to generate an additional chaotic dataset. We use the generator circuit dynamics in Table 1 to generate time series that show quasi-periodic behavior. For each simulated system, we use a single time series corresponding to the $x$ variable (see Fig. 1(a)). In addition, we generate noisy signals of Gaussian white noise, pink noise, and red noise, using the algorithm described by Timmer and Koenig[51]. In the deterministic classes, we produce noisy datasets by adding Gaussian noise to the state variable $x$. The standard deviation of the noise is set as a fixed proportion of the original data, as described below:
@@
[…]
where $\sigma_{x}$ denotes the standard deviation of the data $x(t)$ obtained from the dynamical system, and $\sigma_{n}$ denotes the standard deviation of noise. We normalize each dataset before using it to reconstruct the phase space trajectory and analyze recurrence features, as detailed in the following section.
Table 1: Details of dynamical and stochastic systems used to generate different classes of time series data for training and testing.
@@
Chen 4D [9] $\begin{aligned} \dot{x}&=a(y-x)+e\,yz\\ \dot{y}&=cx-d\,xz+y+u\\ \dot{z}&=xy-bz\\ \dot{w}&=-k\,y\end{aligned}$ a=35,c=25a=35,\;c=25,
d=5,e=35d=5,\;e=35,
k=100,b=7.6k=100,\;b=7.6 Hyperchaotic
k=100,b=4.9k=100,\;b=4.9 Hyperchaotic
k=315,b=4.9k=315,\;b=4.9 Chaotic
k=195,b=4.9k=195,\;b=4.9 Chaotic
k=495,b=4.9k=495,\;b=4.9 Periodic
k=800,b=4.9k=800,\;b=4.9 Quasiperiodic
k=590,b=4.9k=590,\;b=4.9 Quasiperiodic
Lorenz 4D [28] $\begin{aligned} \dot{x}&=a(y-x)\\ \dot{y}&=bx-xz-cy+w\\ \dot{z}&=xy-dz\\ \dot{w}&=-k\,y-r\,w\end{aligned}$ a=12,b=23a=12,\;b=23,
c=1,d=2.1c=1,\;d=2.1,
k=4.00,r=0.20k=4.00,\;r=0.20 Hyperchaotic
k=6.00,r=0.10k=6.00,\;r=0.10 Hyperchaotic
k=5.00,r=0.20k=5.00,\;r=0.20 Hyperchaotic
k=6.00,r=1.00k=6.00,\;r=1.00 Chaotic
k=0.10,r=0.20k=0.10,\;r=0.20 Chaotic
k=0.30,r=0.20k=0.30,\;r=0.20 Chaotic
k=6.00,r=0.45k=6.00,\;r=0.45 Periodic
k=20.0,r=0.20k=20.0,\;r=0.20 Periodic
Lorenz 3D [14] $\begin{aligned} \dot{x}&=\sigma(y-x)\\ \dot{y}&=x(\rho-z)-y\\ \dot{z}&=xy-\beta z\end{aligned}$ $\sigma=10$ $\beta=\tfrac{8}{3}$ $\rho=28$ Chaotic
Generator circuit dynamics [24] $\begin{aligned} \dot{x}&=y\\ \dot{y}&=(\lambda+z+x^{2}-\beta x^{4})y-\omega_{0}^{2}x\\ \dot{z}&=b(\varepsilon-z)-ky^{2}\end{aligned}$ $\lambda=-1$ $\beta=\tfrac{1}{18}$ $\omega_{0}=2\pi$ $b=1$ $k=0.02$ $\varepsilon=4$ Quasiperiodic
Stochastic data [51] $\begin{aligned} N_{1},N_{2}&\sim\mathcal{N}(0,1)\\ S(f_{k})&\sim f_{k}^{-\beta}\\ F(f_{k})&=\sqrt{\tfrac{S(f_{k})}{2}}(N_{1}+iN_{2})\\ x(t)&=\mathcal{F}^{-1}[F(f_{k})]\end{aligned}$ $\beta=0$ White Noise
$\beta=1$ Pink Noise
$\beta=2$ Red Noise
Figure 1: Classification of time series. Schematic representation of the proposed methodology for classifying time series using RPs as input to the proposed deep learning model. (a) We begin by selecting a normalized time series from some simulated dynamical systems and noise generator as listed in Table 1. (b) The phase space of the selected time series is reconstructed using Takens’ embedding theorem. (c) A recurrence plot is then generated from the embedded trajectory using a suitable recurrence threshold $\epsilon$. (d) The resulting RP is resized and used as the input to the deep learning model. In this model, ResNet-50 helps to extract features from RP. Its output is processed through two parallel branches. The standard branch captures high-level spatial features, and the directional branch uses multiple early feature maps from the ResNet 50 pipeline, which are shown by the two arrows drawn from the mid-layers of ResNet 50. The outputs from both branches are concatenated and passed through a dense layer. Finally, a softmax layer generates class wise probabilities, from which the most probable dynamical state is identified. For example, if the input time series is chaotic, the output layer assigns a higher probability to the chaotic class.
2.2 Reconstruction of Phase space
From the normalized univariate time series, we reconstruct the system’s phase space using Takens’ embedding theorem [38]. According to this theorem, the dynamical structure of the original multi-dimensional state space can be recovered by constructing delay vectors with appropriate delay and embedding dimension as,
@@ 2.2 Reconstruction of Phase space
Various techniques are available to estimate appropriate values for the time delay $\tau$ and the embedding dimension $d$, which are essential to reconstruct the trajectory of the system from a single scalar time series. In our study, we employ the mutual information method [49] to estimate $\tau$. The first local minimum of the mutual information curve is typically selected as the optimal delay, as it marks the point at which successive values of the time series become sufficiently independent for effective phase space reconstruction. Although the precise value of the delay may depend on the resolution of the time series, the method remains robust provided that the search is performed over a sufficiently large range to capture the local minimum. To determine the embedding dimension $d$, we use the method of False Nearest Neighbors (FNN), following the approach proposed by Kennel et al. [20]. The underlying idea is that if the reconstructed trajectory is folded onto itself in a lower-dimensional space, it results in a higher number of false neighbors. The knee point at which this number levels off to zero indicates a suitable embedding dimension that preserves the system’s topology.
With the appropriate choice of $\tau$ and $d$ for each time series sample from the systems, the original univariate time series is transformed into a sequence of embedded vectors. These vectors are organized into a matrix $\mathbf{X}$ as follows:
@@
2.3 Recurrence Plots
Recurrence plots provide a two-dimensional representation of the times when a dynamical system revisits states that are close to those encountered previously, within a predefined distance threshold $\epsilon$. Given the delay embedded matrix $\mathbf{X}$ in Eq. 3, the recurrence matrix $\mathbf{R}$ is constructed with elements: